# Removing Reflections from RAW Photos

### v2

@@ F.3 Upsampler comparisons
Figure A18: Upsampling high frequency texture. V-DESIRR [36] re-introduces reflections that were removed at $256^{2}$ pixels. Our method removes these, but textures that were not visible at $256^{2}$ are copied.
In Fig. A17 we provide additional comparison to V-DESIRR [36] in which we upsample a ground truth reflection image (a transmission image is upsampled in Fig. 6). A magnified region is inset, and shows that our method preserves details of a framed photo through the upsampling process, whereas V-DESIRR re-introduces the reflection content (the bars of a fence; see captured photo, white box). In the rest of the image, V-DESIRR introduces strong color artifacts around sharp edges. This is a consistent issue that appears to result from the propagation and amplification of small errors that are made at low resolutions. Our model reduces error propagation by masking $t$ and $r$ out of the high-resolution mixture features at each level. We are able to predict an effective, high resolution mask with a lightweight and fast model by matching features, whereas V-DESIRR must use its model capacity to infer from the high-resolution mixture image what details to add into the low resolution clean images. This latter problem is more difficult to solve with limited model capacity, and it is difficult to avoid propagating errors once they are made.
In Fig. A18 we show a typical failure mode of our model. The input image (top and center left) has a reflection of a textured exterior wall. Our base model correctly removes this reflection at $256^{2}$ pixels (center right). Notice that, at low resolution, it is possible to represent the vertical edges of the exterior wall, but not the texture of the wall. We upsampled this result with both V-DESIRR and our model. V-DESIRR re-introduces the low frequency edges that were removed by the base model, and it copies the high frequency wall texture into the transmission as well. Our model does not re-introduce the low frequency edges, but the high frequency texture of the wall that is not present at low resolution is copied into the transmission image. This produces a texture artifact in the final result. Future work should reduce these kinds of errors while keeping the architecture lightweight. This might be done by reusing feature information from the base model, and across levels of the Gaussian pyramid as the upsampler is iteratively applied.
Figure A19: Predicting the reflection enables aesthetic editing and error correction.
@@ F.4 Editing reflections
In Fig. 7 and Fig. A19 we show examples of reflection editing for aesthetic control and error correction (see also Sec. 5.4). For Fig. A19, a photographer was asked to finish the photo using the outputs of the reflection removal system. They chose to re-introduce the reflection for aesthetic purposes, and to correct errors. The reflections from the edges of the top record player cover were removed by our system (white arrows); the photographer pulled them back into their edited image. Editing was performed in Photoshop using the tone-mapped transmission and reflection images, and the “Linear Dodge (Add)” layer blend mode.
@@ Appendix G Adobe Camera RAW, DNG SDK
Here we detail the code within the DNG SDK [1] version 1.7.1 that is used in our simulation functions Func. 1, Func. A2 and Func. A3. In this work, the necessary SDK functionality was transliterated into Python to interoperate with the geometric simulation (Appendix B) and mixture search (Appendix D). To simplify the exposition, we however describe the SDK code here in a functional manner, whereas the SDK is a class system. Consequently, some of the SDK functions use class member variables, and not all functions in this exposition of the code map one-to-one with functions of the same name in the SDK.
As discussed in Sec. 3 and Appendix A, reflections are simulated in XYZ color space by using the color processing of the DNG SDK, which supports two paths to a white balanced, linear RGB image: the ForwardMatrix or the ColorMatrix. Only the latter path facilitates conversion to a device-independent color space (XYZ) before white balancing, as required for reflection simulation. We therefore implement in Func. A4 the ACR color process in which the ColorMatrix is used (cf. Func. A5). Note that both paths account for the as-shot illuminant because the ColorMatrix is interpolated according to the as-shot illuminant (see Func. A10 and Func. A8).
@@
Function A4 Adobe DNG SDK, convert RAW images to XYZ.
Input: A RAW image
@@
2:  Subtract the Stage-3 black level from I. {SDK Func. A6}
@@
4:  Compute WhiteXY. {SDK Func. A10}
5:  Compute the transform XYZ_to_CAM from WhiteXY. {SDK Func. A8}
6:  Recover saturated highlights in I. {SDK Func. A7}
7:  Transform I to XYZ using inv(XYZ_to_CAM). {see also SDK Func. A5}
@@
    Output: Transform to linear RGB
This function is included here as reference to the entire computation of the color transform to linear RGB. Note, the DNG SDK implements the DNG Spec. [1]. It uses the ForwardMatrix when it is available, and otherwise uses the CameraMatrix. In this work we use only the CameraMatrix, since this supports white balancing after conversion to XYZ.
@@
Function A6 Adobe DNG SDK, get Stage3BlackLevel
@@
    Output: The black level
The Stage-3 black level is not stored in the DNG EXIF header. It is a global scalar offset that remains after spatially varying black levels have been removed by parsing and applying the black level tags. By default, the DNG SDK Stage-3 image has a black level of zero, and negative noise values have been clipped to zero. In this work, we adopt that convention and clip the negative noise for simplicity. Clipping can however be disabled, and the non-zero Stage-3 black level can be recovered with a minor modification to the dng_validate binary, described below. The black level can then be read from the printed output of dng_validate when the stage-3 image is extracted with the -3 option.
@@
Function A7 Adobe DNG SDK, recover saturated highlights
@@
Function A8 Adobe DNG SDK, FindXYZtoCamera
@@
Function A9 Adobe DNG SDK, NeutralToXY (projected, cf. SDK Func. A11)
@@
2:     Compute the XY coordinate of $K$. {SDK Func. A16}
3:     Compute the transform XYZ_to_CAM from XY. {SDK Func. A8}
4:     Compute the XYZ coordinate of XY. {SDK Func. A17}
@@
Function A10 Adobe DNG SDK, computing WhiteXY and CameraWhite
@@
Function A11 Adobe DNG SDK, NeutralToXY
@@
Function A12 Adobe DNG SDK, compute XYZ_to_sRGB
@@
1:  Get the transform XYZ_D50_to_sRGB. {SDK Func. A13}
2:  Get the D50 XY coordinate XY_D50. {SDK Func. A15}
3:  Compute matrix XYZ_to_XYZ_D50 using XY_D50 and XYPoint. {SDK Func. A14}
@@
Function A13 Adobe DNG SDK, get XYZ_D50_to_sRGB
@@
Function A14 Adobe DNG SDK, MapWhiteMatrix
@@
Function A15 Adobe DNG SDK, D50_xy_coord
@@
Function A16 Adobe DNG SDK, Get_xy_coord.
@@
Function A17 Adobe DNG SDK, XYtoXYZ.
@@
Function A18 Adobe DNG SDK, sRGB_to_linear_sRGB.

### v3

@@ F.3 Upsampler comparisons
In Fig. S11 we provide additional comparison to V-DESIRR [35] in which we upsample a ground truth reflection image (a transmission image is upsampled in Fig. 5). A magnified region is inset, and shows that our method preserves details of a framed photo through the upsampling process, whereas V-DESIRR re-introduces the reflection content (the bars of a fence; see captured photo, white box). In the rest of the image, V-DESIRR introduces strong color artifacts around sharp edges. This is a consistent issue that appears to result from the propagation and amplification of small errors that are made at low resolutions. Our model reduces error propagation by masking $t$ and $r$ out of the high-resolution mixture features at each level. We are able to predict an effective, high resolution mask with a lightweight and fast model by matching features, whereas V-DESIRR must use its model capacity to infer from the high-resolution mixture image what details to add into the low resolution clean images. This latter problem is more difficult to solve with limited model capacity, and it is difficult to avoid propagating errors.
In Fig. S14 we show a typical failure mode of our model. The input image (top and center left) has a reflection of a textured exterior wall. Our base model correctly removes this reflection at $256^{2}$ pixels (center right). Notice that, at low resolution, it is possible to represent the vertical edges of the exterior wall, but not the texture of the wall. We upsampled this result with both V-DESIRR and our model.
V-DESIRR re-introduces the low frequency edges that were removed by the base model, and it copies the high frequency wall texture into the transmission as well. Our model does not re-introduce the low frequency edges, but the high frequency texture of the wall that is not present at low resolution is copied into the transmission image. This produces a texture artifact in the final result. Future work should reduce these kinds of errors while keeping the architecture lightweight. This might be done by reusing feature information from the base model, and across levels of the Gaussian pyramid as the upsampler is iteratively applied.
@@ F.4 Editing reflections
In Fig. 6 and Fig. S15 we show examples of reflection editing for aesthetic control and error correction (see also Sec. 5.4). For Fig. S15, a photographer was asked to finish the photo using the outputs of the reflection removal system. They chose to re-introduce the reflection for aesthetic purposes, and to correct errors. The reflections from the edges of the top record player cover were removed by our system (white arrows); the photographer pulled them back into their edited image. Editing was performed in Photoshop using the tone-mapped transmission and reflection images, and the “Linear Dodge (Add)” layer blend mode.
@@ Appendix G Adobe Camera RAW, DNG SDK
Here we detail the code within the DNG SDK [1] version 1.7.1 that is used in our simulation functions Func. 3.3, Func. A.1 and Func. A.1. In this work, the necessary SDK functionality was transliterated into Python to interoperate with the geometric simulation (Appendix B) and mixture search (Appendix D). To simplify the exposition, we however describe the SDK code here in a functional manner, whereas the SDK is a class system. Consequently, some of the SDK functions use class member variables, and not all functions in this exposition of the code map one-to-one with functions of the same name in the SDK.
As discussed in Sec. 3 and Appendix A, reflections are simulated in XYZ color space by using the color processing of the DNG SDK, which supports two paths to a white balanced, linear RGB image: the ForwardMatrix or the ColorMatrix. Only the latter path facilitates conversion to a device-independent color space (XYZ) before white balancing, as required for reflection simulation. We therefore implement in Func. G the ACR color process in which the ColorMatrix is used (cf. Func. G). Note that both paths account for the as-shot illuminant because the ColorMatrix is interpolated according to the as-shot illuminant (see Func. G and Func. G).
@@
Function S3 Adobe DNG SDK, convert RAW images to XYZ.
Input: A RAW imag
@@
2:  Subtract the Stage-3 black level from I. {SDK Func. G}
@@
4:  Compute WhiteXY. {SDK Func. G}
5:  Compute the transform XYZ_to_CAM from WhiteXY. {SDK Func. G}
6:  Recover saturated highlights in I. {SDK Func. G}
7:  Transform I to XYZ using inv(XYZ_to_CAM). {see also SDK Func. G}
@@
Output: Transform to linear RGB This function is included here as reference to the entire computation of the color transform to linear RGB. Note, the DNG SDK implements the DNG Spec. [1]. It uses the ForwardMatrix when it is available, and otherwise uses the CameraMatrix. In this work we use only the CameraMatrix, since this supports white balancing after conversion to XYZ.
@@
Function S5 Adobe DNG SDK, get Stage3BlackLevel
@@
Output: The black level The Stage-3 black level is not stored in the DNG EXIF header. It is a global scalar offset that remains after spatially varying black levels have been removed by parsing and applying the black level tags. By default, the DNG SDK Stage-3 image has a black level of zero, and negative noise values have been clipped to zero. In this work, we adopt that convention and clip the negative noise for simplicity. Clipping can however be disabled, and the non-zero Stage-3 black level can be recovered with a minor modification to the dng_validate binary, described below. The black level can then be read from the printed output of dng_validate when the stage-3 image is extracted with the -3 option.
@@
Function S6 Adobe DNG SDK, recover saturated highlights
@@
Function S7 Adobe DNG SDK, FindXYZtoCamera
@@
Function S8 Adobe DNG SDK, NeutralToXY (projected, cf. SDK Func. G)
@@
2:     Compute the XY coordinate of $K$. {SDK Func. G}
3:     Compute the transform XYZ_to_CAM from XY. {SDK Func. G}
4:     Compute the XYZ coordinate of XY. {SDK Func. G}
@@
Function S9 Adobe DNG SDK, computing WhiteXY and CameraWhite
@@
Function S10 Adobe DNG SDK, NeutralToXY
@@
Function S11 Adobe DNG SDK, compute XYZ_to_sRGB
@@
1:  Get the transform XYZ_D50_to_sRGB. {SDK Func. G}
2:  Get the D50 XY coordinate XY_D50. {SDK Func. G}
3:  Compute matrix XYZ_to_XYZ_D50 using XY_D50 and XYPoint. {SDK Func. G}
@@
Function S12 Adobe DNG SDK, get XYZ_D50_to_sRGB
@@
Function S13 Adobe DNG SDK, MapWhiteMatrix
@@
Function S14 Adobe DNG SDK, D50_xy_coord
@@
Function S15 Adobe DNG SDK, Get_xy_coord.
@@
Function S16 Adobe DNG SDK, XYtoXYZ.
@@
Function S17 Adobe DNG SDK, sRGB_to_linear_sRGB.